40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 9 00:15:21 2009, Yoichi, Update, PSL, Laser head temperature oscillation laser.png
    Reply  Sat May 9 02:20:11 2009, Jenne, Update, PSL, Laser head temperature oscillation 
       Reply  Sat May 9 15:19:10 2009, rana, Update, PSL, Laser head temperature oscillation Picture_2.png
          Reply  Mon May 11 11:49:20 2009, steve, Update, PSL, MOPA cooling water lines are backwards htempvar80d.jpg
             Reply  Thu May 21 18:07:32 2009, rana, Update, PSL, Screw on Needle valve loosened Untitled.png
Message ID: 1569     Entry time: Sat May 9 02:20:11 2009     In reply to: 1568     Reply to this: 1570
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Laser head temperature oscillation 

Quote:
After the laser cooling pipe was unclogged, the laser head temperature has been oscillating in 24h period.
The laser power shows the same oscillation.
Moreover, there is a trend that the temperature is slowly creeping up.
We have to do something to stop this.
Or Rob has to finish his measurements before the laser dies.


How's DTEC doing? I thought DTEC was kind of in charge of dealing with these kinds of things, but after our laser-cooling-"fixing", DTEC has been railed at 0, aka no range.

After glancing at DTEC with Dataviewer along with HTEMP and AMPMON (my internet is too slow to want to post the pic while ssh-ed into nodus), it looks like DTEC is oscillating along with HTEMP in terms of frequency, but perhaps DTEC is running out of range because it is so close to zero? Maybe?
ELOG V3.1.3-